40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Tue Feb 2 18:35:47 2016, gautam, Update, Green Locking, Lightwave NPRO moved from PSL table to SP table 
    Reply  Tue Feb 2 18:54:02 2016, Koji, Update, Green Locking, Lightwave NPRO moved from PSL table to SP table 
       Reply  Wed Feb 3 23:23:47 2016, gautam, Update, Green Locking, Lightwave NPRO moved from PSL table to SP table KnifeEdgeScans_x.pdfBeamscan_x.pdf
          Reply  Thu Feb 4 09:16:46 2016, Koji, Update, Green Locking, Lightwave NPRO moved from PSL table to SP table 
             Reply  Fri Feb 5 00:23:01 2016, gautam, Update, Green Locking, Lightwave NPRO moved from PSL table to SP table Beamscan_setup.pdfBeamscan_x.pdfBeamscan_y.pdfZscan.pdf
                Reply  Fri Feb 5 16:50:24 2016, gautam, Update, Green Locking, First pass at mode-matching Modematch_AUXx.pdfNewSetUp.pngModematch_AUXx_2.pdfXendModeMatch.m.zip
                   Reply  Mon Feb 8 15:36:37 2016, gautam, Update, Green Locking, Alternative mode-matching scheme Modematch_alternative.pdf
                      Reply  Tue Feb 9 04:37:10 2016, ericq, Update, Green Locking, Laser swap initiated 
                         Reply  Tue Feb 9 11:49:47 2016, gautam, Update, Green Locking, Laser swap initiated Innolight_beamProfile.pdf
                            Reply  Tue Feb 9 19:15:36 2016, gautam, Update, Green Locking, Laser swap - updates Modematch_X.pdfXendModeMatch.m.zip
                               Reply  Wed Feb 10 17:57:15 2016, gautam, Update, Green Locking, Laser swap - updates 
                                  Reply  Fri Feb 12 17:05:40 2016, gautam, Update, Green Locking, Laser swap - green light recovered but no flashes in the arm 
                                     Reply  Fri Feb 12 19:07:52 2016, Koji, Update, Green Locking, Laser swap - green light recovered but no flashes in the arm 
                                        Reply  Fri Feb 19 00:42:19 2016, gautam, Update, Green Locking, Laser swap - beam ellipticity from laser? IMG_6287.JPGIMG_6288.JPGIMG_6286.JPG
                                           Reply  Mon Feb 22 13:56:52 2016, gautam, Update, Green Locking, Laser swap -reflected beam from ETM aligned 
                                              Reply  Tue Feb 23 23:01:16 2016, gautam, Update, Green Locking, Laser swap - Green PDH locking AM_scan.pdfAM_scan_zoomed.pdf
                                                 Reply  Wed Feb 24 19:29:13 2016, gautam, Update, Green Locking, Laser swap - Green PDH locked 
                                                    Reply  Fri Feb 26 01:52:44 2016, gautam, Update, Green Locking, Laser swap - Green PDH OLTF X_PDH_OLTF_20160225.pdf
                                                       Reply  Mon Feb 29 17:17:26 2016, gautam, Update, Green Locking, Laser swap - still no green beatnote 
                                                          Reply  Wed Mar 2 17:42:19 2016, gautam, Update, Green Locking, Laser swap - some progress 
                                                             Reply  Thu Mar 3 01:25:50 2016, gautam, Update, Green Locking, Laser swap - 2 IR + 1 green beatnotes found IR_beat_20160303_2.pdf
                                                                Reply  Fri Mar 4 01:11:41 2016, gautam, Update, Green Locking, Laser swap - both green beatnotes found IR_beat_20160303_green.pdf
                                                                   Reply  Sat Mar 5 23:31:01 2016, gautam, Update, Green Locking, Laser swap - some updates AUXxTelescope.png.pngPM_setup.pdfPLLolg.pdfAMPM20160303.pdfIRbeat_20160304.pdf
                                                                      Reply  Mon Mar 7 23:51:36 2016, gautam, Update, Green Locking, Laser swap - some improvement PomonaTF.pdfXPDH.pdfYPDH.pdfgreenbeat_20160307.pdf
                                                                         Reply  Tue Mar 8 18:22:20 2016, rana, Update, Green Locking, Laser swap - some improvement 
Message ID: 11973     Entry time: Wed Feb 3 23:23:47 2016     In reply to: 11971     Reply to this: 11974
Author: gautam 
Type: Update 
Category: Green Locking 
Subject: Lightwave NPRO moved from PSL table to SP table 

As Koji pointed out in the previous elog, the CCD beam profiler was ill suited for this measurement. Nevertheless, to get a rough idea of the beam profile, I made a few rearrangements to my earlier setup:

  • Kept the HWP at the same place it was, as this is roughly the configuration that is going to be used at the endtable anyways. It was ~7cm from the shutter housing on the laser head (unfortunately, I neglected to take a picture).
  • Moved the PBS downstream till it was ~40 cm away (so as to minimize the thermal lensing effect from the ~300mW beam) from the laser head. Rotated the HWP till I got about 6mW of transmitted power (dumped the rest into a black hole)
  • Installed a 95% reflecting BS to further attenuate the power to a level suitable for the CCD (dumped the reflected part onto a razor beam dump)
  • Installed the CCD beam profiler and captured an image, at ~60cm from laser head. In this configuration, I was able to get a clean image capture without the CCD saturating. Unfortunately, I could not transfer it off the laptop used to operate the beam profiler, I will upload a screen capture tomorrow once I get it. Anyways, the main observation was that the beam appeared quite elliptical (ellipticity ~0.6). It was also not clear to what extent thermal lensing at the PBS/BS was afftecting this measurement.

Following Koji's suggestion, I decided to do a knife-edge measurement as well. The measurement configuration was similar to the one described above, except the PBS/BS were removed, and a 1.0 neutral density filter was was installed ~80cm from the laser head (here the ~300 mW beam was >2mm in diameter, as judged by eye). I used the Ophir power meter, which was why I had to install an ND filter as it is rated for 100mW max power. I will put a picture up tomorrow. Thermal lensing shouldn't be of much consequence here, as we just need the whole beam to fall onto the power meter active area (verified by eye), and only the relative change in power levels as the knife edge cuts the beam matters. I took the cross-sectional profile of the beam by translating the knife in the x-direction (i.e. cut the beam "left to right" ).

Attachments 1 and 2 are the results from todays measurements. It remains to repeat by cutting the beam along the y direction, and see what ellipticity (if any) shows up. I also found some "nominal" numbers in page 4 of the Lightwave datasheet - it tells us to expect a waist 5cm from the shutter housing, with horizontal and vertical 1/e^2 diameters of 0.5mm and 0.38mm respectively. My measurement suggests a horizontal diameter of ~0.25mm (half the "nominal" value?!), and the waist location to be 8.22cm from the shutter housing. I wonder if this discrepancy is a red flag? Could it be due to the HWP? I'm reasonably sure of my calculations, and the fits have come out pretty nicely as well...

Attachment 1: KnifeEdgeScans_x.pdf  91 kB  | Show | Show all
Attachment 2: Beamscan_x.pdf  82 kB  | Show | Show all
ELOG V3.1.3-